OpenStudy (kitkat16):

An entrepreneur estimates that the total profit (profit = total revenue minus total cost) from his proposed company will be given by the function P(x) = x^3 – 4x^2 + 3x – 12, where P is in hundreds of dollars and x is the number of years elapsed after the company starts operations. In how many years (x) will the company break even (no profit, no loss)?

OpenStudy (kitkat16):

P(x) = x^3 – 4x^2 + 3x – 12

OpenStudy (anonymous):

Four. \[4^3-4(4^2)+3(4)-12=0\] 4^3= 64 4*4^2=64 3*4=12 64-64+12-12=0

OpenStudy (kitkat16):

so is it 4?

OpenStudy (kitkat16):

lol nevermind I'm so brain dead